A jar contains 5 red marbles numbered 1 to 5 and 8 blue marbles numbered 1 to 8 . A marble is drawn at random from the jar. Find the probability the marble is

a. Even-numbered given that the marble is red.

b. Red given that the marble is even-numbered.


1
Expert's answer
2021-11-09T17:09:48-0500

"P(Red) = \frac{5}{13}"


"P(Blue)=\frac{8}{13}"


"P(Even)=\frac{6}{13}"


"P(odd)=\frac{7}{13}"


P(Even given Red) = "P(Even|Red)=\frac{2}{5}"


P(Red given Even) = "P(Red|Even)=\frac{2}{6}=\frac{1}{3}"




(a) "P(Even|Red)=\frac{P(Red|Even)×P(Even)}{P(Red)}"


"=>P(Even|Red)=\frac{\frac{2}{6}×\frac{6}{13}}{\frac{5}{13}}=\frac{2}{5}"



(b) "P(Red|Even)=\frac{P(Even|Red)×P(Red)}{P(Even)}"


"=>P(Red|Even)=\frac{\frac{2}{5}×\frac{5}{13}}{\frac{6}{13}}=\frac{1}{3}"